OpenStudy (anonymous):

Construct examples a. show that if f and g are one to one, then f+g might not be one-to-one. b. show that if f and g are onto, then f+g need not be onto. c. give an example of functions f and g that are one to one and onto, but f times g is not one to one and not onto.

OpenStudy (zzr0ck3r):

how many times are you going to ask this question?

OpenStudy (perl):

part a) hint: pick a one to one function for f, it can be any. then for g pick its negative. if you add them what do you get?

OpenStudy (perl):

For example f(x) = e^x + 1 g(x) = -e^x both functions are one to one. What happens when you add the functions

OpenStudy (perl):

Actually a simpler function is f(x) = x g(x) = -x f and g are both one to one and onto. However f(x) + g(x) = 0 which is a constant function. The constant function is neither one to one nor onto.

OpenStudy (perl):

f times g is neither one to one nor onto as well

OpenStudy (zzr0ck3r):

@perl the user keeps posting the question under many different names, and refuses to take part